Solve the differential equation
dy/dx + (x/1-x^2)y = x√y, y(0) = 1.
1
Expert's answer
2020-04-20T12:36:26-0400

Given

"y' +\left( \frac{x}{1-x^2}\right)y=xy^{1/2}"

This is a Bernoli equation , multiply both sides by "y^{-1\/2}" then

"y^{-1/2}y' +\left( \frac{x}{1-x^2}\right)y^{1/2}=x"

Let "z=y^{1/2},\ \ \ \ \ ,z'= \frac{1}{2}y^{-1/2}y'" , then

"2z' +\left( \frac{x}{1-x^2}\right)z=x\\
z' +\frac{1}{2}\left( \frac{x}{1-x^2}\right)z=\frac{x}{2}"

This is a linear equation with integral coefficient

"\mu= e^{\int\left( \frac{ x}{2(1-x^2)}\right)dx }\\
\ \ \ \ = e^{-\frac{1}{4}\ln (1-x^2)}\\
\ \ \ \ =(1-x^2)^{-1/4}"

Then the general solution is

"\mu z \ \ \ \ \ \ \ \ = \frac{1}{2} \int \mu (x)dx\\
(1-x^2)^{-1/4}z= \frac{1}{2} \int x(1-x^2)^{-1/4}dx\\
(1-x^2)^{-1/4}z=-\frac{1}{3}\left(1-x^2\right)^{\frac{3}{4}}+C\\
 (1-x^2)^{-1/4}\sqrt{y}=-\frac{1}{3}\left(1-x^2\right)^{\frac{3}{4}}+C"

Since "y(0)=1\ \ \ \to C= \frac{4}{3}"

"(1-x^2)^{-1/4}\sqrt{y}=-\frac{1}{3}\left(1-x^2\right)^{\frac{3}{4}}+\frac{4}{3}\\
\ \ \ \ \ \ \ \ \ \ \ \sqrt{y}= \frac{-1}{3}(1-x^2) +\frac{4}{3}(1-x^2)^{1/4}\\
\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ y=\left[ \frac{-1}{3}(1-x^2) +\frac{4}{3}(1-x^2)^{1/4}\right]^2"